# A Systematic Literature Review on Detecting Software Vulnerabilities with Large Language Models

### v1

@@ 4.4.3. Learning Paradigms
Adversarial training enhances model robustness by introducing crafted perturbed code examples during training, improving resilience to noise and attacks (Yu et al., 2022; Chen et al., 2025; Huynh et al., 2025).
Sankey diagram showing the relationship between system architecture, input representation, adaptation technique, and task formulation in surveyed studies.
Figure 2 maps the relationships between system architecture, input representation, adaptation techniques, and task formulation across the 227 surveyed studies. Each node is annotated with the absolute number of studies assigned to that category, providing a quantitative overview of the landscape. The diagram enables a holistic view of how LLMs are combined with the different input representations and adaptation strategies. The rightmost part of the diagram visualizes how these techniques correlate with task formulations, highlighting the dominance of binary classification. The diagram further reveals underexplored combinations, such as the use of structure-aware inputs with PEFT and prompting methods for vulnerability-specific classification. This overview serves as a compact visual summary of the field and offers starting points for researchers to explore novel system designs or revisit rarely studied combinations.
@@ 5. Vulnerability Datasets
In this SLR, we set an emphasis on investigating the datasets used in the training and evaluation of LLM-based software vulnerability detection approaches. Datasets play a central role in the development and evaluation: Their quality, structure, and coverage of diverse vulnerability types directly affect model performance and generalization. Dataset choice can influence which methods appear to perform better, potentially favoring certain vulnerability types. Systematically comparing datasets is essential to understand their suitability for the vulnerability detection task, assess their limitations, and ensure meaningful evaluations.
To systematically analyze the datasets used in the surveyed studies, we extend the presented taxonomy to cover datasets. We outline the taxonomy based on the datasets used most commonly for training or evaluation in the surveyed studies, discussing type, granularity, source, and labeling. Further, we investigate the CWE coverage and diversity of selected datasets, discussing limitations such as long-tail distributions and lack of representative vulnerability coverage. Finally, we discuss trends in dataset use, including comparability of evaluations across different use cases.
@@ 5.1. Taxonomy and Commonly Used Datasets
We categorize vulnerability datasets according to type, granularity, source, and labelling, as shown in Figure 3, to capture their quality and realistic nature. Similar to related surveys (Shereen et al., 2024; Ferrag et al., 2025a; Taghavi Far and Feyzi, 2025; Sheng et al., 2025; Guo et al., 2024a; Guo and Bettaieb, 2023), we provide a comprehensive overview of commonly used datasets in the surveyed studies as well as selected, more recent datasets suited to address open challenges of LLM-based methods, refer to Table 3. We briefly introduce these datasets along with the taxonomy and discuss the resulting practical challenges and biases in dataset construction. In addition, we compare the selected datasets based on further characteristics such as their programming languages covered, as well as quantitative characteristics, i.e., dataset size, the share of samples representing vulnerabilities, and the number of distinct CWEs covered, cf. Table 4, as a proxy for label diversity and comprehensiveness.
Taxonomy showing categories and attribute values for vulnerability detection datasets used.
@@ 5.1.1. Dataset Type
Real-World datasets are sourced from open-source software projects, e.g., Devign (Zhou et al., 2019), or curated vulnerability databases, e.g., Big-Vul (Fan et al., 2020) and MegaVul (Ni et al., 2024b). These datasets are more suited for a realistic evaluation, capturing the challenges posed by detecting software vulnerabilities in real-world projects. However, due to the efforts required for crawling and parsing software projects, real-world datasets often suffer from label noise.
@@ 5.1.2. Dataset Granularity
Medium-grained datasets segment code at the Program, Slice or Function level. The largest share of commonly used vulnerability datasets is labeled at the function level (Metrics and Evaluation, 2006; Russell et al., 2018; Zhou et al., 2019; Fan et al., 2020; Zheng et al., 2021; Chakraborty et al., 2022; Bhandari et al., 2021; Chen et al., 2023; Ding et al., 2024a; Du et al., 2025; Ni et al., 2024b; Li et al., 2025f), cf. Table 3. The VulDeePecker dataset, also known as the Code Gadget Database (CGD), represents program slices as code gadgets composed of a number of program statements that are semantically related to each other through data or control dependencies. While this granularity supports a more focused analysis, it may miss critical contextual information when vulnerabilities span multiple functions or require interprocedural analysis.
Fine-grained datasets label vulnerabilities at the Line level enable precise vulnerability localization. Big-Vul, D2A (Zheng et al., 2021), FormAI, and ReposVul provide annotations at the line-level. Such a narrow scope, however, omits the broader context required to detect certain vulnerabilities, especially those involving complex data flows or control dependencies.
@@
| SARD (Metrics and Evaluation, 2006) | 2006 | Mixed | File | Open-Source | Synthetically | 28 | (Metrics and Evaluation, 2006) |
@@
| VulDeePecker (Li et al., 2018) | 2018 | Mixed | Slice | Constructed | Security Vendor | 9 | (CGCL Codes, 2017) |
| Draper (Russell et al., 2018) | 2018 | Mixed | Function | Constructed | Tool, Synthetically | 8 | (Kim and Russell, 2018) |
| Devign (Zhou et al., 2019) | 2019 | Real | Function | Open-Source | Developer | 53 | (Pinconschi, 2020) |
| Big-Vul (Fan et al., 2020) | 2020 | Real | Function, Line | Collected | Security Vendor | 50 | (Fan, 2020) |
| D2A (Zheng et al., 2021) | 2021 | Real | Function, Line | Open-Source | Tool | 10 | (IBM, 2021) |
| ReVeal (Chakraborty et al., 2022) | 2021 | Real | Function | Collected | Developer | 30 | (VulDetProject, 2020) |
| CVEfixes (v1.0.8) (Bhandari et al., 2021) | 2021 | Real | Function, Commit, File | Collected | Security Vendor | 14 | (Moonen and Vidziunas, 2024) |
@@
| SecurityEval (v2.1) (Siddiq and Santos, 2022) | 2022 | Mixed | Program | Constructed | Synthetically, Tool | 5 | (Security & Software Engineering Research Lab at University of Notre Dame, 2022) |
| DiverseVul (Chen et al., 2023) | 2023 | Real | Function | Collected | Developer | 21 | (Wagner Research Group, 2025) |
| SVEN (He and Vechev, 2023) | 2023 | Real | Program | Constructed | Developer | 4 | (SRI Lab, ETH Zurich, 2023) |
| FormAI* (Tihanyi et al., 2023) | 2023 | Synthetic | Program, Line | Constructed | Tool | 1 | (FormAI Dataset, 2023) |
| ReposVul* (Wang et al., 2024a) | 2024 | Real | Project, File, Function, Line | Collected | Tool | 2 | (Wang, 2024) |
| PrimeVul (Ding et al., 2025) | 2024 | Real | Function | Constructed | Security Vendor | 15 | (DLVulDet, 2024) |
| PairVul* (Du et al., 2025) | 2024 | Real | Function | Collected | Security Vendor | 1 | (KnowledgeRAG4LLMVulD, 2024) |
| MegaVul* (2024/04) (Ni et al., 2024b) | 2024 | Real | Function | Collected | Security Vendor | 2 | (Icyrockton, 2023) |
| CleanVul* (Li et al., 2025f) | 2024 | Real | Function | Open-Source | Developer, Tool | 1 | (Li, 2024) |
@@ 5.1.3. Dataset Source
Open-Source datasets are derived from public data collections accessible through open-source platforms or repositories, e.g., Github repositories. Examples for open-source datasets are SARD, Devign, D2A, and CleanVul (Li et al., 2025f). Their accessibility ensures transparency and reproducibility, making them widely used in academic research. However, datasets sourced directly from Github repositories often lack reliable ground-truth labels and are sensitive to repository selection and curation quality, i.e., the specific projects chosen and how samples are filtered, labeled, and balanced, leading to biased, noisy, or unrepresentative data.
@@
Closed-Source datasets are obtained from commercial or industrial entities. Closed-sourced datasets are absent from the listed commonly used datasets. Notably, only two surveyed studies (Ding et al., 2023; Xue et al., 2025) use an industrial, closed-sourced dataset. Although such datasets offer significant potential for research targeting real-world deployment scenarios, they are often subject to restrictions that limit the publication and sharing of company-internal data. Such restricted access limits reproducibility, public benchmarking, and broader adoption within the research community.
@@ 5.1.4. Dataset Labeling
Tool-created labels are automatically generated using, e.g., static analyzers or formal verification tools. For instance, in the Draper dataset, labels are generated using static analyzers and then manually mapped to binary vulnerability labels and CWEs by a team of security experts. Similarly, the D2A dataset applies differential analysis to VFC version pairs from several open-source projects using static analysis tools, identifying disappearing bugs likely to present true vulnerabilities. Notably, the FormAI dataset (Tihanyi et al., 2023) consists of compilable and self-contained C programs generated using GPT-3.5 turbo, incorporating varying levels of complexity to simulate a broad spectrum of coding patterns. Labels are derived using formal verification. ReposVul jointly employs LLMs for evaluating the relevance between code changes and vulnerability fixes, and static analysis tools for checking the vulnerabilities in the changes. While tools enable large-scale dataset creation, they are often criticized for generating a high volume of false positives. Static analysis tools, in particular, produce numerous warnings, many of which originate from low-confidence detections (Croft et al., 2023). Formal verification offers higher detection reliability but at a higher computational cost.
@@
| SARD (Metrics and Evaluation, 2006) | C/C++, C#, Java, PHP | File | n/s | ✓ | 150+ | |
| Juliet C/C++(Metrics and Evaluation, 2017a) | C/C++ | File | n/s | ✓ | 118 | |
| Juliet Java (Metrics and Evaluation, 2017b) | Java | File | n/s | ✓ | 112 | |
| VulDeePecker (Li et al., 2018) | C/C++ | Slice | 28.7 | ✓ | 2 | |
| Draper (Russell et al., 2018) | C/C++ | Function | 6.5 | ✓ | 4 | |
| Devign (Zhou et al., 2019) | C/C++ | Function | 48.0 | ✗ | n/a | |
| Big-Vul (Fan et al., 2020) | C/C++ | Function, Line | 4.5 | ✓ | 91 | |
| D2A (Zheng et al., 2021) | C/C++ | Function, Line | 1.4 | ✓ | n/a | |
| ReVeal (Chakraborty et al., 2022) | C/C++ | Function | 9.2 | ✗ | n/a | |
| CVEfixes (v1.0.8) (Bhandari et al., 2021) | 27 languages | Function, Commit, File | 45.5 | ✓ | 272 | |
| CrossVul (Nikitopoulos et al., 2021b) | 40+ languages | File | 50.0 | ✓ | 168 | |
| SecurityEval (v2.1) (Siddiq and Santos, 2022) | Python | Program | 100.0 | ✓ | 69 | |
| DiverseVul (Chen et al., 2023) | C/C++ | Function | 5.7 | ✓ | 150 | |
| SVEN (He and Vechev, 2023) | C/C++, Python | Program | 50.0 | ✓ | 9 | |
| FormAI* (Tihanyi et al., 2023) | C | Program, Line | 80.2 | (✓) | (41) | |
| ReposVul* (Wang et al., 2024a) | C/C++, Java, Python | Project, File, Function, Line | 0.7 | ✓ | 236 | |
| PrimeVul (Ding et al., 2025) | C/C++ | Function | 2.9 | ✓ | 140 | |
| PairVul* (Du et al., 2025) | C/C++ | Function | 50.0 | ✓ | 95 | |
| MegaVul* (2024/04) (Ni et al., 2024b) | C/C++ | Function | 5.1 | ✓ | 176 | |
| CleanVul* (Li et al., 2025f) | C/C++, C#, Java, JavaScript, Python | Function | 50.0 | ✗ | n/a |

### v2

@@ 4.4.2. Adaptation Technique (T2)
Adversarial training enhances model robustness by introducing crafted perturbed code examples during training, improving robustness to noise and attacks (Yu et al., 2022; Chen et al., 2025b; Zhao and Guan, 2025; Du et al., 2024; Liu et al., 2026).
4.4.3. Orchestration Technique (T3)
While adaptation techniques focus on optimizing the model’s internal parameters or input context, orchestration techniques focus on the design of the inference process. Rather than relying on a simple input-model-output pipeline, orchestration techniques embed the LLM into advanced system designs that utilize multiple generation steps, iterative refinement, or collaborative workflows. We categorize orchestration techniques into Multi-Step, Verification, Agentic, Ensemble, and Controller-based systems.
Some studies engage the model in a Multi-Step (T3.1) analysis, rather than issuing a single prompt for vulnerability detection (Espinha Gasiba et al., 2024; Huynh et al., 2025; Liu et al., 2024f, 2023; Ullah et al., 2024; Wang et al., 2025b; Yu et al., 2025; Jiao et al., 2025; Yang et al., 2025b; Mhatre et al., 2025; Peng et al., 2025; Tsai et al., 2025a; Ding et al., 2025). A common strategy within this approach is to decouple comprehension from detection, prompting the model to first analyze the code’s functionality or intent before querying for vulnerabilities (Espinha Gasiba et al., 2024; Huynh et al., 2025; Liu et al., 2024f, 2023; Jiao et al., 2025; Yang et al., 2025b; Peng et al., 2025; Wang et al., 2025b). Beyond functional analysis, Yu et al. (Yu et al., 2025) task the model to first generate supplementary code files based on commit messages, subsequently using this supplementary context for the final detection. Mhatre et al. (Mhatre et al., 2025) follow a debugging style, first identifying code issues, followed by a deeper, targeted analysis of specific issues. Tsai et al. (Tsai et al., 2025a) propose a sequential approach that dynamically selects the optimal strategy, choosing between direct zero-shot classification, RAG-based prediction, or multi-agent collaboration, guided by the model’s internal confidence and collaborative signals.
To mitigate hallucinations and improve precision, studies increasingly adopt Verification (T3.2) mechanisms where models are prompted to review and improve generated outputs. Verification extends the standard prompt-response interaction by re-prompting the model to evaluate, validate, or refine a prediction. In single-model setups, the model is typically prompted to assess its prior response by verifying the correctness of its output (Xia et al., 2025) or by recursively critiquing and improving (Tamberg and Bahsi, 2025; Du et al., 2024; Sayagh and Ghafari, 2025). In multi-model settings, verification is implemented as an interactive process between models (collaborative refinement), possibly under different roles. For example, using the output of one model to prompt another model with corrective cues such as ”An expert has found… please recheck it” (Wang et al., 2025d; Ambati et al., 2024; Tsai et al., 2025b; Widyasari et al., 2024; Cao and Jun, 2024; Eberhardt and Milánkovich, 2024).
Agentic (T3.3) systems embed LLMs into autonomous, decision-making loops, following a higher-level behavioral pattern involving multi-step planning, tool use, and self-directed reasoning. Agents may operate individually or as part of multi-agent systems (Cao and Jun, 2024; Yildiz et al., 2025; Widyasari et al., 2025; Ahmed et al., 2025; Eberhardt and Milánkovich, 2024; Li et al., 2025h; Tsai et al., 2025a; Wang et al., 2025e; Tsai et al., 2025b), where agents specialize in subtasks or roles, interact with tools, and collaboratively refine outputs. For example, Yildiz et al. (Yildiz et al., 2025) employ Reasoning and Acting (ReAct) agents (Yao et al., 2023), which follow a thought-action-observation cycle, for vulnerability detection in code repositories. Widyasari et al. (Widyasari et al., 2025) present a courtroom-inspired multi-agent framework with four role-specific agents: security researcher, code author, moderator, and review board. The agents collaborate to assess code vulnerabilities through iterative analysis, argumentation, summarization, and a final verdict based on collective reasoning. Ahmed et al. (Ahmed et al., 2025) propose a multi-agent pipeline consisting of a normalization, planning, context, detection, and validation agent. Sayagh and Ghafari (Sayagh and Ghafari, 2025) use multiple agentic teams, with a lister searching for potential CWEs in the input code, a reviewer requesting missing CWEs, further agents identifying external code context, and a final agent deciding on each CWE based on gathered context.
Ensemble (T3.4) approaches combine the representations or predictions of diverse models to mitigate individual model limitations and biases. To leverage complementary feature representations, e.g., semantic, contextual, and syntactic features, studies fuse representations from different pre-trained models (Du et al., 2025b; Wu and Xiao, 2025; Mahmud et al., 2025; Ridoy et al., 2024; Shimmi et al., 2024a; Xiong and Dong, 2024; Sun et al., 2023; JianJie and Le, 2023). For example, Ridoy et al. (Ridoy et al., 2024) employ an ensemble stacking approach to synthesize the semantic understanding of CodeBERT, the structural representations of GraphCodeBERT, and the cross-modal capabilities of UniXcoder. Distinct from standard feature fusion, Tian et al. (Tian et al., 2025) introduce a parameter fusion method that merges independent classification heads trained on distinct vulnerability datasets, allowing the model to adapt to new vulnerability types. Ensemble approaches further include decision-level aggregation via voting mechanisms to combine predictions from multiple LLMs (Widyasari et al., 2024; Zhang et al., 2025b) or ensembles of LLMs and other tools (Zhang et al., 2025b; Zhou et al., 2024a).
Moving beyond static inference pipelines, some studies use LLMs in a Controller (T3.5) or collaboration setting. A central model or routing mechanism analyzes the input code and coordinates downstream resources, selecting, e.g., the most appropriate expert models or prompts. Yang et al. (Yang et al., 2025a) and Wu and Xiao (Wu and Xiao, 2025) implement a Mixture-of-Experts (MoE) framework where a router dynamically assigns input code to the most appropriate vulnerability experts. Similarly, Wang et al. (Wang et al., 2025e) employ a meta-agent to route code to specialized analyzers. Peng et al. (Peng et al., 2025) introduce a Mixture-of-Prompts architecture that activates specific prompt experts (dynamic prompt pipelines) based on code features of the code under detection. Li et al. (Li et al., 2025c) simulate team collaboration involving a leader and various expert models specialized on syntactic features; the leader orchestrates the process and derives the final detection from the experts’ analyses. Farr et al. (Farr et al., 2025) extend the routing logic to human collaborators. They flag code as automatic quarantine, cleared for deployment, or human analyst evaluation, where code with high model uncertainty is routed to human experts.
@@ 5. Vulnerability Datasets
In this SLR, we set an emphasis on investigating the datasets used in the adaptation and evaluation of LLM-based vulnerability detection approaches. Datasets play a central role: Their quality, structure, and coverage of diverse vulnerability types directly affect model performance and generalization. Further, the choice of dataset does not merely benchmark performance. A critical challenge in this domain is data leakage, where data samples overlap with the corpora used to pre-train the model, potentially leading to inflated performance.
Furthermore, the use of datasets has evolved beyond standard training. With the rise of advanced prompting techniques, datasets now serve as a dynamic input context, providing exemplars for in-context prompting or knowledge retrieval. This shift also changes which parts of the data are used. While deep learning-based models often relied solely on code syntax, LLMs actively leverage the natural language components of datasets, such as commit messages, vulnerability descriptions, and code comments, to support high-level explanation and reasoning tasks. Therefore, systematically comparing datasets is essential to understand their suitability for LLM-based detection, assess their limitations, and ensure meaningful evaluations.
To systematically analyze the datasets used in the surveyed studies, we extend the presented taxonomy to cover datasets. We outline the taxonomy based on the datasets used most commonly for training or evaluation in the surveyed studies, discussing type, granularity, source, and labeling. Further, we investigate the CWE coverage and diversity of selected datasets, discussing limitations such as long-tail distributions and a lack of representative vulnerability coverage. Finally, we discuss trends in dataset usage, emphasizing the comparability of evaluations across different use cases.
@@ 5.1. Taxonomy and Commonly Used Datasets
We categorize vulnerability datasets according to type, granularity, source, and labeling, as shown in Figure 6. Similar to related surveys (Shereen et al., 2024; Ferrag et al., 2025a; Taghavi Far and Feyzi, 2025; Sheng et al., 2025; Guo et al., 2024a; Guo and Bettaieb, 2023), we provide a comprehensive overview of commonly used datasets in the surveyed studies as well as selected, more recent datasets suited to address open challenges of LLM-based approaches, refer to Table 4. We briefly introduce these datasets along with the taxonomy and discuss the resulting practical challenges and biases in dataset construction. In addition, we compare the selected datasets based on further characteristics such as their programming languages covered, as well as quantitative characteristics, i.e., dataset size, the share of samples representing vulnerabilities, and the number of distinct CWEs covered (cf. Table 5) as a proxy for label diversity and comprehensiveness.
@@ 5.1.1. Dataset Type
Real-World datasets are sourced from open-source software projects, e.g., Devign (Zhou et al., 2019), or curated vulnerability databases, e.g., Big-Vul (Fan et al., 2020) and MegaVul (Ni et al., 2024). These datasets are more suited for a realistic evaluation, capturing the challenges posed by detecting software vulnerabilities in real-world projects. However, due to the efforts required for crawling and parsing software projects, real-world datasets often suffer from label noise.
@@ 5.1.2. Dataset Granularity
Medium-grained datasets segment code at the Program, Slice or Function level. The largest share of commonly used vulnerability datasets is labeled at the function level (Metrics and Evaluation, 2006; Russell et al., 2018; Zhou et al., 2019; Fan et al., 2020; Zheng et al., 2021; Chakraborty et al., 2022; Bhandari et al., 2021; Chen et al., 2023; Ding et al., 2024a; Du et al., 2025a; Ni et al., 2024; Li et al., 2025k), cf. Table 4. The VulDeePecker dataset, also known as the Code Gadget Database (CGD), represents program slices as code gadgets composed of a number of program statements that are semantically related to each other through data or control dependencies. While this granularity supports a more focused analysis, it may miss critical contextual information when vulnerabilities span multiple functions or require interprocedural analysis.
Fine-grained datasets label vulnerabilities at the Line level, enabling precise vulnerability localization. Big-Vul, D2A (Zheng et al., 2021), FormAI, and ReposVul provide annotations at the line-level. Such a narrow scope, however, omits the broader context required to detect certain vulnerabilities, especially those involving complex data flows or control dependencies.
@@
| SARD (Metrics and Evaluation, 2006) | 2006 | Mixed | File | Open-Source | Synthetically | 32 | (Metrics and Evaluation, 2006) |
@@
| VulDeePecker (Li et al., 2018) | 2018 | Mixed | Slice | Constructed | Security Vendor | 11 | (CGCL Codes, 2017) |
| Draper (Russell et al., 2018) | 2018 | Mixed | Function | Constructed | Tool, Synthetically | 11 | (Kim and Russell, 2018) |
| Devign (Zhou et al., 2019) | 2019 | Real | Function | Open-Source | Developer | 69 | (Pinconschi, 2020) |
| Big-Vul (Fan et al., 2020) | 2020 | Real | Function, Line | Collected | Security Vendor | 63 | (Fan, 2020) |
| D2A (Zheng et al., 2021) | 2021 | Real | Function, Line | Open-Source | Tool | 9 | (IBM, 2021) |
| ReVeal (Chakraborty et al., 2022) | 2021 | Real | Function | Collected | Developer | 40 | (VulDetProject, 2020) |
| CVEfixes (v1.0.8) (Bhandari et al., 2021) | 2021 | Real | Function, Commit, File | Collected | Security Vendor | 19 | (Moonen and Vidziunas, 2024) |
@@
| SecurityEval (v2.1) (Siddiq and Santos, 2022) | 2022 | Mixed | Program | Constructed | Synthetically, Tool | 4 | (Security & Software Engineering Research Lab at University of Notre Dame, 2022) |
| DiverseVul (Chen et al., 2023) | 2023 | Real | Function | Collected | Developer | 27 | (Wagner Research Group, 2025) |
| SVEN (He and Vechev, 2023) | 2023 | Real | Program | Constructed | Developer | 8 | (SRI Lab, ETH Zurich, 2023) |
| FormAI* (Tihanyi et al., 2023) | 2023 | Synthetic | Program, Line | Constructed | Tool | 2 | (FormAI Dataset, 2023) |
| ReposVul* (Wang et al., 2024a) | 2024 | Real | Project, File, Function, Line | Collected | Tool | 3 | (Wang, 2024) |
| PrimeVul (Ding et al., 2025) | 2024 | Real | Function | Constructed | Security Vendor | 27 | (DLVulDet, 2024) |
| PairVul* (Du et al., 2025a) | 2024 | Real | Function | Collected | Security Vendor | 1 | (KnowledgeRAG4LLMVulD, 2024) |
| MegaVul* (2024/04) (Ni et al., 2024) | 2024 | Real | Function | Collected | Security Vendor | 2 | (Icyrockton, 2023) |
| CleanVul* (Li et al., 2025k) | 2024 | Real | Function | Open-Source | Developer, Tool | 3 | (Li, 2024) |
@@ 5.1.3. Dataset Source
Open-Source datasets are derived from public data collections accessible through open-source platforms or repositories, e.g., Github repositories. Examples for open-source datasets are SARD, Devign, D2A, and CleanVul (Li et al., 2025k). Their accessibility ensures transparency and reproducibility, making them widely used in academic research. However, datasets sourced directly from GitHub repositories often lack reliable ground-truth labels and are sensitive to repository selection and curation quality, i.e., the specific projects chosen and how samples are filtered, labeled, and balanced, leading to biased, noisy, or unrepresentative data.
@@
Closed-Source datasets are obtained from commercial or industrial entities. Closed-sourced datasets are absent from the listed commonly used datasets. Notably, only few surveyed studies (Ding et al., 2023; Mock et al., 2025; Peng et al., 2025; Xue et al., 2025) use an industrial, closed-sourced dataset. Although such datasets offer significant potential for research targeting real-world deployment scenarios, they are often subject to restrictions that limit the publication and sharing of company-internal data. Such restricted access limits reproducibility, public benchmarking, and broader adoption within the research community.
@@ 5.1.4. Dataset Labeling
Tool-created labels are automatically generated using, e.g., static analyzers or formal verification tools. For instance, in the Draper dataset, labels are generated using static analyzers and then manually mapped to binary vulnerability labels and CWEs by a team of security experts. Similarly, the D2A dataset applies differential analysis to VFC version pairs from several open-source projects using static analysis tools, identifying disappearing bugs likely to present true vulnerabilities. Notably, the FormAI dataset (Tihanyi et al., 2023) consists of compilable and self-contained C programs generated using GPT-3.5-Turbo, incorporating varying levels of complexity to simulate a broad spectrum of coding patterns. Labels are derived using formal verification. ReposVul jointly employs LLMs for evaluating the relevance between code changes and vulnerability fixes and static analysis tools for checking the vulnerabilities in the changes. While tools enable large-scale dataset creation, they are often criticized for generating a high volume of false positives. Static analysis tools, in particular, produce numerous warnings, many of which originate from low-confidence detections (Croft et al., 2023). Formal verification offers higher detection reliability but at higher computational costs.
@@
| SARD (Metrics and Evaluation, 2006) | C/C++, C#, Java, PHP | File | 450,000+ | n/s | ✓ | 150+ |
| Juliet C/C++(Metrics and Evaluation, 2017a) | C/C++ | File | 64,099 | n/s | ✓ | 118 |
| Juliet Java (Metrics and Evaluation, 2017b) | Java | File | 28,881 | n/s | ✓ | 112 |
| VulDeePecker (Li et al., 2018) | C/C++ | Slice | 61,638 | 28.7 | ✓ | 2 |
| Draper (Russell et al., 2018) | C/C++ | Function | 1,274,366 | 6.5 | ✓ | 4 |
| Devign (Zhou et al., 2019) | C/C++ | Function | 48,687 | 48.0 | ✗ | n/a |
| Big-Vul (Fan et al., 2020) | C/C++ | Function, Line | 264,919 | 4.5 | ✓ | 91 |
| D2A (Zheng et al., 2021) | C/C++ | Function, Line | 1,295,623 | 1.4 | ✓ | n/a |
| ReVeal (Chakraborty et al., 2022) | C/C++ | Function | 18,169 | 9.2 | ✗ | n/a |
| CVEfixes (v1.0.8) (Bhandari et al., 2021) | 27 languages | Function, Commit, File | 277,948 | 45.5 | ✓ | 272 |
| CrossVul (Nikitopoulos et al., 2021b) | 40+ languages | File | 27,476 | 50.0 | ✓ | 168 |
| SecurityEval (v2.1) (Siddiq and Santos, 2022) | Python | Program | 121 | 100.0 | ✓ | 69 |
| DiverseVul (Chen et al., 2023) | C/C++ | Function | 330,492 | 5.7 | ✓ | 150 |
| SVEN (He and Vechev, 2023) | C/C++, Python | Program | 1,606 | 50.0 | ✓ | 9 |
| FormAI* (Tihanyi et al., 2023) | C | Program, Line | 246,549 | 80.2 | (✓) | (41) |
| ReposVul* (Wang et al., 2024a) | C/C++, Java, Python | Project, File, Function, Line | 232,465 | 0.7 | ✓ | 236 |
| PrimeVul (Ding et al., 2025) | C/C++ | Function | 235,768 | 2.9 | ✓ | 140 |
| PairVul* (Du et al., 2025a) | C/C++ | Function | 8,628 | 50.0 | ✓ | 95 |
| MegaVul* (2024/04) (Ni et al., 2024) | C/C++ | Function | 353,873 | 5.1 | ✓ | 176 |
| CleanVul* (Li et al., 2025k) | C/C++, C#, Java, JavaScript, Python | Function | 11,632 | 50.0 | ✗ | n/a |